# Quantifying the Dunkelflaute: An analysis of variable renewable energy droughts in Europe

### v1

@@ Abstract
Variable renewable energy droughts, also referred to as “Dunkelflaute”, emerge as a challenge for realizing climate-neutral energy systems based on variable wind and solar power. Using data on 38 historic weather years and an advanced identification method, we characterize European drought events for on- and offshore wind power, solar photovoltaics, and policy-relevant renewable technology portfolios. We show that drought characteristics heavily depend on the chosen threshold. Using single thresholds, as common in the literature, is thus not advisable. Applying a multi-threshold framework, we quantify how the complementarity of wind and solar power temporally and spatially alleviates drought frequency, duration, and severity within (portfolio effect) and across countries (balancing effect). We further identify the most extreme droughts and show how these drive major discharging periods of long-duration storage in a fully renewable European energy system. Such events comprise sequences of shorter, contiguous droughts of varying severity. In a perfectly interconnected Europe, the most extreme drought event occurred in winter 1996/97 and lasted 55 days. Yet, the average renewable portfolio availability during this event was still 47% of its long-run mean. As extreme droughts may span across the turn of years, single calendar year planning horizons are not suitable for modeling weather-resilient future energy scenarios.
@@ 1 Introduction
Replacing fossil fuels with renewable energy sources is a major strategy for achieving climate neutrality. However, the potential for dispatchable renewable energy sources such as hydro power, geothermal, or bioenergy is limited in most countries. Hence, future renewable energy systems will heavily rely on variable wind and solar power [iea_net_2023]. With increasing penetration of these variable renewable energy (VRE) sources, energy systems become increasingly exposed to weather variability. Spatial or temporal system flexibility is needed to match variable renewable energy supply with demand. This includes geographical balancing via transmission, different types of energy storage, and demand response [denholm_grid_2011, rasmussen_storage_2012, schlachtberger_benefits_2017, roth_geographical_2023, he_sector_2021].
@@
Dealing with extreme periods of low wind and solar availability emerges as a key challenge for realizing energy systems with very high shares of VRE sources. Often referred to as “Dunkelflauten”, such VRE drought periods are characterized by long-lasting and substantial shortages of VRE supply [kittel_measuring_2024] and may cover large geographical areas, limiting the potential for spatial smoothing across affected countries. Dealing with VRE drought events necessitates the use of long-duration storage and other flexibility options [dowling_role_2020]. Hence, understanding the spatial and temporal characteristics of VRE drought events is crucial for weather-resilient energy system planning and energy policy as well as for the design of energy markets and support instruments for generation and flexibility technologies. This includes questions as to how frequent, how long, and how severe such periods are, and assessing their spatial and temporal correlation across large-scale interconnected energy systems.
@@
Literature from different fields contributes to VRE drought analysis. Wind droughts in the UK are well-studied [cannon_using_2015, potisomporn_spatial_2022, potisomporn_evaluating_2023, abdelaziz_assessing_2024], focusing on frequency-duration distributions, return periods as well as spatial and temporal correlations of historic or future on- and offshore wind droughts. Similar analyses for wind power have been conducted for Ireland [leahy_persistence_2013], the North Sea [patlakas_low_2017], Germany [ohlendorf_frequency_2020], or, analyzing deviations from climatological means, globally [antonini_identification_2024]. Additionally, research interest in drought patterns of policy-relevant portfolios comprising wind and solar photovoltaics (PV) is growing. Historic, future, and synthetic weather data of various world regions have been analyzed, such as Europe [raynaud_energy_2018, kies_critical_2021, kapica_potential_2024, hu_implications_2023, breyer_reflecting_2022], the U.S. [rinaldi_wind_2021], India [gangopadhyay_role_2022], Germany [kaspar_climatological_2019, mockert_meteorological_2023], Hungary [mayer_probabilistic_2023], and Japan [ohba_climatology_2022]. Besides regional and seasonal variations for single technologies and VRE portfolios, a general finding is that combining wind and solar within regions, as well as considering balancing across regions, can mitigate drought characteristics.
@@
Various methodological approaches have been used for renewable drought identification [ohlendorf_frequency_2020, potisomporn_extreme_2024, kittel_measuring_2024]. Droughts can be defined as periods of consecutive time steps with renewable availability below a certain drought threshold, either with fixed [raynaud_energy_2018, rinaldi_wind_2021, kies_critical_2021, gangopadhyay_role_2022, kapica_potential_2024, kaspar_climatological_2019, ohba_climatology_2022] or variable duration [potisomporn_extreme_2024], using various identification methods. For instance, these periods can be identified by searching for an availability constantly below the threshold [potisomporn_spatial_2022, potisomporn_evaluating_2023, abdelaziz_assessing_2024, ohlendorf_frequency_2020, kies_critical_2021, breyer_reflecting_2022, kaspar_climatological_2019, mayer_probabilistic_2023, ohba_climatology_2022], a mean availability over a certain averaging interval below the threshold [abdelaziz_assessing_2024, leahy_persistence_2013, patlakas_low_2017, ohlendorf_frequency_2020, raynaud_energy_2018, rinaldi_wind_2021, gangopadhyay_role_2022, kapica_potential_2024, mockert_meteorological_2023], or the deviation of the availability from a drought threshold [potisomporn_extreme_2024] or its climatological mean [antonini_identification_2024, stoop_climatological_2024]. Drought thresholds are either exogenously and presumably arbitrarily set [cannon_using_2015, potisomporn_extreme_2024, abdelaziz_assessing_2024, leahy_persistence_2013, patlakas_low_2017, ohlendorf_frequency_2020, kies_critical_2021, kaspar_climatological_2019, mockert_meteorological_2023, mayer_probabilistic_2023, ohba_climatology_2022], or derived from the data analyzed, such as a fraction of the time-invariant mean [raynaud_energy_2018, rinaldi_wind_2021, kapica_potential_2024] or maximum [breyer_reflecting_2022] availability, or of its time-variant climatological mean [gangopadhyay_role_2022, antonini_identification_2024, stoop_climatological_2024]. As each of these approaches has strengths and weaknesses, no standard drought identification method has yet emerged in the literature [kittel_measuring_2024].
@@
Here we analyze and compare VRE drought events for single renewable technologies as well as policy-relevant VRE portfolios for 33 European countries (EU27, the UK, Norway, Switzerland, and the Western Balkans) and for a pan-European “copperplate” scenario. In the latter case, we assume perfect interconnection across all countries that allows for unconstrained geographical balancing. We draw on historic VRE availability factor time series covering 38 years [de_felice_entso-e_2022]. We use an advanced, open-source algorithm that fully captures unique drought periods, properly accounts for brief periods of higher renewable availability within longer drought events, and avoids arbitrary threshold choices [kittel_measuring_2024]. We provide evidence on how frequent, how long, and how severe such periods have been in the past, their seasonality, and to what extent they are correlated across the interconnected European energy system in space and time.
@@
Our results quantify how the spatial and temporal complementarity of wind and solar PV alleviates the issue of VRE droughts within European countries (“portfolio effect”). Averaged over all thresholds and countries, the maximum drought duration of a renewable technology portfolio that combines solar PV as well as on- and offshore wind power decreases by 64%, 52%, or 47% compared to standalone PV, onshore wind, or offshore wind droughts. Our results also demonstrate how European integration can further mitigate VRE droughts by leveraging spatially complementary VRE availability profiles across regions (“balancing effect”). Considering unconstrained geographical balancing across all countries, the longest technology portfolio drought shortens by 65%. We further show that drought characteristics strongly depend on the chosen drought threshold and that single-threshold analyses lead to an incomplete characterization of extreme drought patterns. Using a multi-threshold analysis, we illustrate that the most extreme VRE events manifest as sequential droughts with varying severity, affecting multiple countries simultaneously to different extents. Based on this, we introduce a drought mass indicator to identify the most extreme droughts and illustrate how these determine long-duration storage needs in a fully renewable European energy system. In Germany, the most extreme event occurred in the winter of 1995/96 and lasted 109 days. Under the assumption of unconstrained geographical balancing across Europe, the most extreme storage-defining event was substantially shorter and lasted 55 days, occurring in the winter of 1996/97. Such extreme droughts may occur at the turn of years, suggesting that planning horizons based on single calendar years are inappropriate for modeling weather-resilient future energy systems.
@@ 2 Results
In the main part of the paper, we illustrate selected results for a limited set of countries. The Supplementary Information SI provides additional results. Interactive, high-resolution versions of all graphs as well as an additional animated graph are available online [kittel_high_2024].
@@ 2.1 Frequency-duration distribution: short drought events occur much more often than longer ones
A cumulative frequency-duration distribution shows how often VRE droughts that lasted at least a certain duration occurred on average per year in the investigated data (Figure 1). This visualization is particularly useful to illustrate the occurrence of shorter-duration events, which are relevant for the operation of short-duration flexibility options in the power sector, e.g. battery storage. Note that the longest drought events, which may drive the need for long-duration electricity storage, are not visible here as the axis is truncated. The following sections focus more on such extreme periods. For a given threshold, the frequencies of all events that are at least as long as a given duration are shown in ascending order. Thresholds are scaled relative to the multi-annual mean of the respective time series (see Section 4).
Figure 1: Exemplary cumulative frequency-duration distributions of drought events, sorting the frequencies of all events that are at least as long as a given duration. White space indicates the absence of droughts for given thresholds in the data. The contour lines represent the threshold-specific frequency. For illustration, the distributions are truncated at 360 hours, i.e., they show events with a maximum duration of just above two weeks. Figure SI.2 shows similar distributions for events lasting up to one full year. Frequencies of events lasting at least two days, one week, and a fortnight are marked for a relative threshold of $\tau=0.75$, i.e., the 75%-fraction of the mean availability factor.
@@
The cumulative frequency-duration distribution substantially differs between wind and solar power. For PV and low thresholds, we identify many events that last less than one day. They reflect the typical diurnal fluctuations of solar PV, with zero production at night-time [ueckerdt_analyzing_2015]. Increasing thresholds lead to fewer PV drought events, as many night-time periods of zero electricity generation merge into multi-day events. When the threshold approaches $\tau=1$, the identified droughts increasingly reflect a strong solar seasonality in Europe, with lower availability in winter than in summer. For on- and offshore wind power, droughts lasting less than one day occur less often. This is because wind power does not have regular non-availability at night-time as solar PV, but fluctuates less regularly. In turn, events that last longer than one day are more frequent, reflecting a typical multi-day variability of wind power [apt_spectrum_2007, ueckerdt_analyzing_2015]. The frequency-duration distribution of VRE portfolio droughts reveals a portfolio effect, i.e., droughts are less frequent when wind and solar PV are combined due to complementary wind and solar availability. Hardly any VRE portfolio droughts arise for lower thresholds. For higher thresholds, their frequency still is much lower than for single technologies (compare marked exemplary frequencies in each column in Figure 1).
@@
Droughts occur less often in an ideal European interconnection than in single countries due to a geographical balancing effect. This is true for both single VRE technologies and portfolios. The reason for this is that periods of low solar and wind availability are not perfectly correlated across Europe [roth_geographical_2023]. For example, assuming a threshold $\tau=0.75$, there have been on average approximately eight, five, or four VRE portfolio droughts per year in Germany, Spain, or Europe that lasted at least one week, respectively.
@@
In the SI, we provide additional results on the seasonality of drought patterns for Europe under the assumption of perfect interconnection. For low thresholds and events lasting longer than a few days, PV droughts are more frequent in winter, while wind droughts are more frequent in summer (Figure SI.3). The SI also shows return periods, which are the reciprocals of yearly frequencies. Similar to frequencies, return periods vary substantially between technologies and regions (see Section SI.3.)
@@
2.2 Maximum drought duration strongly depends on threshold
Figure 2: Exemplary most extreme duration of single drought events for each year and across all thresholds. The contour lines represent threshold-specific maximum duration. White space indicates the absence of droughts for given thresholds in the data. The events with the highest and lowest duration across all years are marked for a threshold $\tau=0.75$. Arrows indicate values that are hidden in valleys of the distribution plane.

### v2

@@ Abstract
Variable renewable energy droughts, also called “Dunkelflaute”, emerge as a challenge for climate-neutral energy systems based on variable renewables. Drawing on 38 historic weather years and an advanced identification method, we characterize European drought events for on- and offshore wind power, solar photovoltaics, and renewable technology portfolios. We show that their characteristics heavily depend on the chosen drought threshold, questioning the usefulness of single-threshold analyses. Applying a multi-threshold framework, we quantify how the complementarity of wind and solar power temporally and spatially alleviates drought frequency, duration, and severity within (portfolio effect) and across countries (balancing effect). We identify the most extreme droughts and show how these drive major discharging periods of long-duration storage in a fully renewable European energy system, based on a policy-relevant decarbonization scenario. Such events comprise sequences of shorter droughts of varying severity. The most extreme event occurred in winter 1996/97 and lasted 55 days in a perfectly interconnected setting. While the average renewable availability during this period was still 47% of its long-run mean, we argue that system planners must consider such events when planning for storage and other flexibility technologies. Methodologically, we conclude that using single calendar years is not suitable for modeling weather-resilient energy scenarios.
@@ 1 Introduction
Mitigating climate change requires deep reductions of greenhouse gas emissions that originate from fossil fuel use [lee_ipcc_2023]. Nuclear energy or fossil fuels in combination with carbon capture and storage are options for this [sepulveda_role_2018], but these firm low-carbon technologies face economic challenges and political controversies in many countries. Another major strategy for achieving climate neutrality is shifting energy supply to renewable energy sources, which are globally on the rise as they become increasingly cost-competitive [irena_renewable_2025]. However, the potential for dispatchable renewable energy sources such as hydro power, geothermal, or bioenergy is limited in most countries. Hence, future renewable energy systems are likely to heavily rely on variable wind and solar power [iea_net_2023, way_empirically_2022]. As these variable renewable energy (VRE) sources may become the most relevant primary energy sources in many countries [jacobson_100_2017, brown_response_2018, child_flexible_2019, pickering_diversity_2022, wang_accelerating_2023], energy systems are increasingly exposed to weather variability. In turn, spatial and temporal system flexibility is increasingly needed to match the supply of VRE sources with demand. This includes geographical balancing via transmission, different types of energy storage, and demand response [denholm_grid_2011, rasmussen_storage_2012, schlachtberger_benefits_2017, roth_geographical_2023, he_sector_2021, hunter_techno-economic_2021, tong_geophysical_2021].
@@
While future energy systems with high VRE shares sources have to deal with variability on different timescales, extreme periods of low wind and solar availability emerge as a key challenge for realizing such renewable energy systems. Often referred to as “Dunkelflauten”, these VRE drought periods are characterized by long-lasting and substantial shortages of wind and solar energy [raynaud_energy_2018, kittel_measuring_2024] and may cover large geographical areas.If the availability of firm low-carbon generation technologies is low, dealing with VRE drought events necessitates the use of long-duration storage and other flexibility options [dowling_role_2020, sepulveda_design_2021]. Hence, understanding the spatial and temporal characteristics of VRE drought events is crucial for weather-resilient energy system planning and energy policy as well as for the design of energy markets and support instruments for generation and flexibility technologies. This includes questions as to how frequent, how long, and how severe such periods are, and assessing their spatial and temporal correlation across large-scale interconnected energy systems.
@@
Literature from different fields contributes to VRE drought analysis. Wind droughts in the United Kingdom are well-studied [cannon_using_2015, potisomporn_spatial_2022, potisomporn_evaluating_2023, abdelaziz_assessing_2024], focusing on frequency-duration distributions, return periods as well as spatial and temporal correlations of historic or future on- and offshore wind droughts. Similar analyses for wind power have been conducted for Ireland [leahy_persistence_2013], the North Sea [patlakas_low_2017], Germany [ohlendorf_frequency_2020], or, analyzing deviations from climatological means, globally [antonini_identification_2024]. Additionally, research interest in drought patterns of policy-relevant portfolios comprising wind and solar photovoltaics (PV) is growing. Historic, future, and synthetic weather data of various world regions have been analyzed, such as Europe [raynaud_energy_2018, kies_critical_2021, kapica_potential_2024, hu_implications_2023, breyer_reflecting_2022, van_der_most_temporally_2024], the U.S. [rinaldi_wind_2021], China [li_renewable_2024], India [gangopadhyay_role_2022], Germany [kaspar_climatological_2019, mockert_meteorological_2023], Hungary [mayer_probabilistic_2023], Japan [ohba_climatology_2022], and Australia [richardson_climate_2023]. Besides regional and seasonal variations for single technologies and VRE portfolios, a general finding is that combining wind and solar within regions, as well as considering balancing across regions, can mitigate drought characteristics.
@@
Different methodological approaches have been used for renewable drought identification [ohlendorf_frequency_2020, potisomporn_extreme_2024, kittel_measuring_2024, li_renewable_2024]. Droughts can be defined as periods of consecutive time steps with renewable availability below a certain drought threshold, either with fixed [raynaud_energy_2018, rinaldi_wind_2021, kies_critical_2021, gangopadhyay_role_2022, kapica_potential_2024, kaspar_climatological_2019, ohba_climatology_2022] or variable duration [potisomporn_extreme_2024], using various identification methods. For instance, these periods can be identified by searching for an availability constantly below the threshold [potisomporn_spatial_2022, potisomporn_evaluating_2023, abdelaziz_assessing_2024, ohlendorf_frequency_2020, kies_critical_2021, breyer_reflecting_2022, kaspar_climatological_2019, mayer_probabilistic_2023, ohba_climatology_2022], a mean availability over a certain averaging interval below the threshold [abdelaziz_assessing_2024, leahy_persistence_2013, patlakas_low_2017, ohlendorf_frequency_2020, raynaud_energy_2018, rinaldi_wind_2021, gangopadhyay_role_2022, kapica_potential_2024, mockert_meteorological_2023], or the deviation of the availability from a drought threshold [potisomporn_extreme_2024] or its climatological mean [antonini_identification_2024, stoop_climatological_2024]. Drought thresholds are either exogenously and presumably arbitrarily set [cannon_using_2015, potisomporn_extreme_2024, abdelaziz_assessing_2024, leahy_persistence_2013, patlakas_low_2017, ohlendorf_frequency_2020, kies_critical_2021, kaspar_climatological_2019, mockert_meteorological_2023, mayer_probabilistic_2023, ohba_climatology_2022], or derived from the data analyzed, such as a fraction of the time-invariant mean [raynaud_energy_2018, rinaldi_wind_2021, kapica_potential_2024] or maximum [breyer_reflecting_2022] availability, or of its time-variant climatological mean [gangopadhyay_role_2022, antonini_identification_2024, stoop_climatological_2024]. As each of these approaches has strengths and weaknesses, no standard drought identification method has yet emerged in the literature [kittel_measuring_2024].
@@
Here we analyze and compare VRE drought events for single renewable technologies and for a fully renewable European electricity system, using VRE portfolio assumptions from a policy-relevant decarbonization scenario from European transmission grid operators. We do so for 33 individual European countries (EU27, the United Kingdom, Norway, Switzerland, and the Western Balkans), and for a pan-European “copperplate” scenario with perfect interconnection across all countries that allows for unconstrained geographical balancing. We draw on a large dataset of historical VRE availability factor time series covering 38 years [de_felice_entso-e_2022]. We use an advanced, open-source algorithm that fully captures unique drought periods, properly accounts for brief periods of higher renewable availability within longer drought events, and avoids arbitrary threshold choices [kittel_measuring_2024]. Our analysis illustrates the variety and seasonality of historical renewable energy drought patterns in Europe and provides evidence on how frequent, how long, and how severe such periods have been in the past, and to what extent they are correlated across the interconnected European energy system in space and time.
@@
Our results quantify how the spatial and temporal complementarity of wind and solar PV alleviates VRE droughts within European countries, giving rise to a technology portfolio effect. Averaged over all thresholds and countries, the maximum drought duration of a renewable technology portfolio that combines solar PV as well as on- and offshore wind power decreases by 64%, 52%, or 47% compared to standalone PV, onshore wind, or offshore wind droughts. Our results also demonstrate how European integration can further mitigate VRE droughts by leveraging spatially complementary VRE availability profiles across regions, signified by a balancing effect. Considering unconstrained geographical balancing across all countries, the longest technology portfolio drought shortens by 65%. We further show that drought characteristics strongly depend on the chosen drought threshold and that single-threshold analyses lead to an incomplete characterization of extreme drought patterns. Using a multi-threshold analysis, we illustrate that the most extreme VRE events manifest as sequential droughts with varying severity, affecting multiple countries simultaneously to different extents. We introduce a drought mass indicator to identify the most extreme droughts and illustrate how these determine long-duration storage needs in a fully renewable European energy system. In Germany, the most extreme event occurred in the winter of 1995/96 and lasted 109 days. Under the assumption of unconstrained geographical balancing across Europe, the most extreme storage-defining event was substantially shorter and lasted 55 days, occurring in the winter of 1996/97. Such extreme droughts may occur at the turn of years, suggesting that planning horizons based on single calendar years are inappropriate for modeling weather-resilient future energy systems.
@@ 2 Results
We define a renewable energy drought as a period during which the moving average of the hourly VRE availability factor, also referred to as capacity factor, is below a given qualification threshold [kittel_measuring_2024]. As there is no consensus in the literature on appropriate drought thresholds, we apply a large number of thresholds to each investigated region and technology, ranging in 5% increments from 10% to 100% of the mean renewable availability across all investigated years [kittel_measuring_2024]. While low thresholds identify severe droughts, higher thresholds increasingly include more moderate drought events. By scaling drought thresholds relative to the mean availability factor of all hours in the data of each region-technology setting (henceforth referred to as “relative thresholds”), we account for regional and technology-specific differences in generation potentials. This enables meaningful cross-regional and cross-technological comparisons (for more details, see Section 4.1). To give an example, an identified wind power drought with a 24-hour duration at a threshold $\tau=0.1$ should be interpreted such that less than 10% of the long-run average electricity generation of wind power is available on average during a 24-hour period.
@@
In the main part of the paper, we illustrate selected results for a limited set of countries. The Supplementary Information SI provides additional results. Interactive, high-resolution versions of all graphs that allow for retrieving concrete values for each of the illustrated 3D-plots as well as an additional animated graph are available online [kittel_high_2025].
2.1 Complex renewable energy drought dynamics and complementarities across regions and technologies
Figure 1 shows renewable energy droughts that last longer than one day for the years 1996 and 1997, which emerge to be particularly critical (see Section 2.5). The Figure combines these events for all employed relative thresholds, for selected countries from southern, central and northern Europe, as well as for the pan-European copperplate scenario. These countries represent some of the largest electricity markets in Europe, each with distinct yet complementary wind and solar generation profiles. The pan-European copperplate scenario combines respective profiles from all 33 countries in our data set, assuming unconstrained geographical balancing. The multi-threshold illustration shows the variety of renewable drought patterns in Europe across regions and technologies even within the same time period. It further indicates that the identified drought duration is highly sensitive to the underlying color-coded threshold. In general, single-threshold analyses using low thresholds can detect extreme short-duration droughts, which may occur in isolation or adjacent to high-availability periods (white spaces in the Figure). Higher thresholds increasingly identify longer-lasting events. These potentially include consecutive shorter periods of extreme renewable shortage or surplus, which are smoothed by the averaging mechanics of the identification algorithm (see Section 4.1). The multi-threshold perspective demonstrates that longer-lasting events identified by higher thresholds may encompass multiple severe droughts detected at lower thresholds.
Figure 1: Identified drought patterns in 1996 and 1997 for all investigated relative thresholds $\tau\in[0.1,1]$ and selected regions. For each technology-specific panel, a horizontal band indicates drought occurrences for the color-coded threshold of one region. To illustrate persistent patterns, only droughts lasting longer than one day are displayed. Figure SI.2 focuses on longer-lasting events by illustrating respective patterns for droughts lasting at least one week.
Solar PV events identified by moderate or high relative thresholds can last several months, indicating low solar availability during winter in Europe (Figure 1, panel b). Lower thresholds identify very severe PV droughts driven by low seasonal availability in combination with factors such as persistent cloud or snow coverage, such as in winter 1996. In summer, only rare and isolated PV droughts of lower severity occur that are longer than one day. These seasonal PV characteristics can generally be found across all years and are often more pronounced in Northern Europe, e.g., observable in winter 2011/12 (compare Germany or United Kingdom to Spain in Figure SI.3).
In contrast, severe long-lasting onshore and offshore wind droughts can occur throughout the year but tend to be more frequent in summer (Figure 1, panels c and d), which is in line with general wind seasonality [kaspar_climatological_2019]. The multi-threshold perspective further reveals that severe wind droughts identified by lower thresholds can be sequential and may occur within contiguous below-average wind periods detected at higher thresholds. The latter may last up to several months. While such long-lasting events usually occur in summer (compare Figures SI.3-SI.5), they may also take place in winter in some countries (e.g., in 1996/97, Figure 1). On- and offshore droughts are generally correlated, but may differ in severity (Figure 1, panels c and d).
The multi-threshold illustration also reveals a technology portfolio effect due to complementary wind and solar power droughts. Typically, high solar availability offsets wind droughts in summer, while wind power often compensates for the low solar PV availability in winter. This results not only in briefer but also less severe renewable portfolio droughts, i.e., higher thresholds apply in the same periods compared to single technologies (compare panel a of Figure 1 with panels b-d). However, if severe and persistent wind and PV droughts coincide, pronounced long-lasting compound droughts can occur even for renewable portfolios, as seen for instance in winter 1996/97.
The Figure also illustrates a geographical balancing effect. Assuming perfect interconnection across Europe, geographical balancing can substantially mitigate the severity and duration of renewable energy droughts. This can be seen when comparing the European copperplate scenario (top row in each panel) to the corresponding isolated-countries cases in Figure 1. This balancing effect is particularly pronounced for onshore wind, as onshore wind droughts tend to occur at different times across space, which allows for spatial smoothing across all 33 considered countries. The effect is less pronounced for offshore wind, which can only be deployed in countries with sea access. This limits the spatial extent over which balancing is possible. It is even less pronounced for solar PV because PV droughts are heavily driven by solar seasonality, which affects most countries simultaneously.
Geographical balancing also reduces the severity and duration of technology portfolio drought events as solar seasonality is lower in Southern Europe and wind droughts do not occur simultaneously across countries. Severe and long-lasting portfolio droughts may arise in single countries and also across Europe at the beginning of a calendar year, (e.g., observable in 2012, Figure SI.3), at the end of a year (e.g., 1982, Figure SI.4), or across the turn of years (e.g., in winter 1996/97, Figure 1). In some years, winter portfolio droughts are much less pronounced (e.g., 2013/14, Figure SI.5). Countries relying heavily on wind power, such as the United Kingdom, may also experience very severe events in summer.
While this section highlights important technology-specific and regional characteristics of renewable energy droughts in Europe drawing on specific weather years, it necessarily remains selective. In the following, we analyze the occurrence of European renewable droughts in a more systematic and exhaustive way for all technologies, regions, and weather years in our data, focusing on both shorter and very long-lasting events. In doing so, we also connect to and extend analyses based on single thresholds which have been prevalent in the literature so far.
@@ 2.2 Frequency-duration distribution: short drought events occur much more often than longer ones
A cumulative frequency-duration distribution shows how often VRE droughts that lasted at least a certain duration occurred on average per year in the investigated data. Figure 2 illustrates these distributions across all investigated relative thresholds for Germany and Spain, as examples of central and southern European countries with different wind and solar energy resources, and for the European copperplate scenario. For a given threshold, the frequencies of all events that are at least as long as a given duration are shown in ascending order on the vertical axis. For instance, at a relative threshold $\tau=0.75$, Germany experienced on average 9.8 onshore wind droughts lasting at least 48 hours (two days), 5.2 lasting at least 168 hours (one week), and 3.1 lasting at least 336 hours (two weeks) per year. Note that the duration axis in Figure 2 is truncated at 360 hours, so it focuses on the occurrence of shorter-duration events. These are particularly relevant for the operation of shorter-duration flexibility options in the power sector, e.g., battery or hydro reservoir storage. Longer-lasting drought exceeding 360 hours, which may drive the need for long-duration electricity storage or other firm low-carbon technologies, are not visible. The following sections focus more on such extreme periods.
Figure 2: Example of cumulative frequency-duration distributions of drought events across all investigated relative drought thresholds $\tau\in[0.1,1]$, sorting the yearly frequencies of all events that are at least as long as a given duration. White space indicates the absence of droughts for given thresholds in the data. The contour lines represent the threshold-specific yearly frequency. For illustration, the distributions are truncated at 360 hours, i.e., they show events with a maximum duration of just above two weeks. Figure SI.6 shows similar distributions for events lasting up to one full year. Yearly frequencies of events lasting at least two days, one week, and a fortnight are marked for a relative threshold of $\tau=0.75$.
@@
The cumulative frequency-duration distribution substantially differs between wind and solar power. For PV and low thresholds, we identify many events that last less than one day. They reflect the typical diurnal fluctuations of solar PV, with zero production at night-time [ueckerdt_analyzing_2015]. Increasing thresholds lead to fewer PV drought events, as many night-time periods of zero electricity generation merge into multi-day events. When the threshold approaches $\tau=1$, the identified droughts increasingly reflect a strong solar seasonality in Europe, with lower availability in winter than in summer. For on- and offshore wind power, droughts lasting less than one day occur less often. This is because wind power does not have regular non-availability at night-time as solar PV, but fluctuates less regularly. In turn, events that last longer than one day are more frequent, reflecting a typical multi-day variability of wind power [apt_spectrum_2007, ueckerdt_analyzing_2015].
The frequency-duration distribution of VRE portfolio droughts also illustrates the portfolio effect, i.e., droughts identified by lower thresholds (see white spaces in the portfolio row in Figure 2 are less frequent when wind and solar PV are combined due to complementary wind and solar availability. Hardly any VRE portfolio droughts arise for lower thresholds. For higher thresholds, their frequency still is much lower than for single technologies (compare highlighted frequencies in each column in Figure 2).
@@
The Figure also illustrates the geographical balancing effect, as droughts occur less often in an ideal European interconnection than in single countries. This is true for both single VRE technologies and portfolios. The reason for this is that periods of low solar and wind availability are not perfectly correlated across Europe [roth_geographical_2023]. For example, assuming a relative threshold $\tau=0.75$, there have been on average approximately eight and five VRE portfolio droughts per year in Germany and Spain that lasted at least one week, respectively, while less than four of such events occurred in Europe.
@@
In the SI, we provide additional results on the seasonality of drought patterns for Europe under the assumption of perfect interconnection. For low thresholds and events lasting longer than a few days, PV droughts are more frequent in winter, while wind droughts are more frequent in summer (Figure SI.7). The SI also shows return periods, which are the reciprocals of yearly frequencies. Similar to frequencies, return periods vary substantially between technologies and regions (see Section SI.4.)
@@
2.3 Maximum drought duration strongly depends on threshold and differs across years and seasons